dy dx 2x-1 Q1) find for the following functions a) y = b) y = x3e-* c) y = cos x + d) x3 + y2 = 1 x-3 1 tanx
Answer #1

{ d = vody-undy y your U2 a>. Y = 22-1 a-3. dy (x-3)(2) - (22-1)(0) (x-3)2 21-6-271+ (2-3)2 da = -5 (4-3) 2 Y = x3ex. dy du =